OpenStudy (anonymous):

Please help me I have 2 questions please do not help me if you do not know what you are doing MEDALS AND FANS QUESTIONS BELOW

OpenStudy (anonymous):

@BioEpic @blackops2luvr @mathmath333 please help me

OpenStudy (mathmath333):

 \(\large\tt \begin{align} \color{black}{\text{ perimetre(p)=a+b+c+d}\\~\\ =60\cfrac{5}{6}+54\cfrac{1}{3}+59\cfrac{1}{3}+56\cfrac{1}{6}\\~\\ =\dfrac{60\times 6+5}{6}+\dfrac{54\times 3+1}{3}+\dfrac{59\times 3+1}{3}+\dfrac{56\times 6+1}{6}}\end{align}\)